OpenStudy (anonymous):

the perimeter of a rectangle equals the sum of the lengths of its four sides. If the width of a rectangle is 8 inches and its perimeter is 50 inches, find the length of the rectangle. the length of the rectangle is ?inches

OpenStudy (anonymous):

@Atkinsoha

OpenStudy (anonymous):

@physicsboffin

OpenStudy (anonymous):

16+2x=50

OpenStudy (anonymous):

x is the length

OpenStudy (anonymous):

so @DoYourHomework thts the final answer 16+2=50

OpenStudy (anonymous):

subtract 16 from 50 then divide that answer by 2

OpenStudy (anonymous):

50-16=34 divide that by 2 to get the length

OpenStudy (anonymous):

What @Atkinsoha said

OpenStudy (anonymous):

17

OpenStudy (anonymous):

correct! :)

OpenStudy (anonymous):

thanks 4 being my math tutors for 2day i appreciate it @Atkinsoha and @DoYourHomework

OpenStudy (anonymous):

no problem!

OpenStudy (anonymous):

What @Atkinsoha said